# Space science & the space economy

### v1

@@ 2.3 Costs: satellite production
The second major cost of a space mission is the cost related to the satellite procurement, integration, and test. Table 4 reports the cost of several scientific satellites, including payloads, bridging the typical cost of large flagship missions and small explorer missions (1.4-0.42M\$/kg). As a comparison, we also report the costs of standard cubesats using COTS components, advanced cubesat, cubesats for commercial constellations, and the SpaceX Starlink satellites.
@@
Satellite Multiplicity Weight kg Cost M\$ Cost M\$/kg
@@
The cost/kg of standard cubesats is 5-20 times lower than the cost/kg of Explorer missions. Of course, standard cubesat are very simple and limited. A better comparison can be done with advanced cubesat, designed specifically for a scientific mission. One good example is HERMES Pathfinder (HERMES-PF), a European constellation of 6 3U cubesat equipped with miniaturized but advanced X-ray and gamma-ray detectors to observe cosmic explosions such as Gamma Ray Bursts and the electromagnetic counterparts of gravitational wave events [15, 16]. The net cost of each HERMES-PF 3U CubeSat is $\sim$\$860.000, including payload (P/L). The spacecraft (S/C) uses selected COTS components, procured through contracts with companies including non-recurring engineering phases to tailor the hardware to the specific application. The P/L and the S/C do not use radhard Electrical, Electronic and Electro-mechanical (EEE) components. HERMES-PF S/C cost/kg is about 145.000 \$/kg (the cost/kg of the P/L is similar), or 2-7 times higher than standard 3U cubesat but only one-third of that of Explorer missions. The cost of cubesats and mini/midi-sats for commercial constellations is similar or lower to that of standard cubesats, at a level of 20.000-30.000\$/kg  [^20] https://sustainability.e-shape.eu/economics-behind-satellite-megaconstellations/https://www.bloomberg.com/news/features/2017-06-29/the-tiny-satellites-ushering-in-the-new-space-revolution
The cost of SpaceX Starlink satellites (bus and P/L included) is lower than that of cubesats by more than another order of magnitude, reaching the incredibly low value of about 1000\$/kg [^21] https://spacenews.com/starlink-soars-spacexs-satellite-internet-surprises-analysts-with-6-6-billion-revenue-projection/. Interestingly, this cost/kg is about constant for all Starlink satellites despite a weight difference of a factor of 6. This is 20-30 times lower than the cost of a standard cubesat and that of advanced cubesat and mini-sat for Earth observation, more than 100 times lower than the cost of advanced cubesats for science applications, and more than 400 times lower than the cost of Explorer missions.
@@ 3 Public Funding Pros and Cons for Large Missions
So far science missions in space have been paid almost entirely by public funding (governments) and thus they belong to the ’old space’ economy. Table 5 summarizes pros and cons of this situation.
@@
Figure 2 shows the evolution of NASA and ESA budgets since 2000. NASA’s budget for science, including solar system exploration, Mars exploration, astrophysics, and Earth observation, varied between 25% and 35% of the total between 2000 and 2025. The ESA budget for science decreased from about 15% of the total in 2006-2008 to 8.5% in 2020-2025. This does not include Earth observation, which includes in the ESA system both science and application (the Copernicus program). Figure 2 also shows NASA and ESA science budgets in absolute values, corrected to inflation to 2023. Note the peaks in the NASA budget around 2003-2006 and 2018-2022. The second is mainly due to spending on JWST. The first is due to spending on several smaller missions, including Messenger, MRO, New Horizon, Dawn, Stereo and Fermi, witnessing a period of great diversification in NASA science missions. ESA science budget remained remarkably constant around a value of 600-700 M\$(2023) per year for about 25 years (the 2003 budget was increased by about 100MEUR to take stock of the grounding of Ariane 5, which postponed the launch of Rosetta). ESA budget for Earth observation increased consistently from 2008 to 2025 from 500 M\$(2023) to 2700 M\$ (2023), due to the investment of the European Commission in the Copernicus program.
@@
    Use of space qualified parts. As an example, the RAD750 computer that operates in many NASA missions including the Perseverance Mars rover costs several hundred thousands dollars, and is based on 1990 technology with a PowerPC 130-200MHz processor, $\sim$260-400 DMIPS [^22] RAD750® radiation-hardened PowerPC microprocessor BAE Systems https://www.baesystems.com/en-media/uploadFile/20210404045936/1434555668211.pdf. As a comparison, the computer used by the Ingenuity helicopter on board Perseverance is based on the Qualcomm Flight board, which uses a smartphone computer chip, the Qualcomm Snapdragon 801 processor with a 2.5GHz clock, some 20 times more powerful and 20 times cheaper than the RAD750232323https://www.qualcomm.com/news/onq/2021/03/journey-mars-how-our-collaboration-jet-propulsion-laboratory-fostered-innovation. The Snapdragon 801 worked flawlessly for the about five years of cruise and operations of Ingenuity.
@@
    Failure is not an Option leads to large numbers of quality control procedures to reduce the risk of failure. As an example, the ECSS (European Cooperation for Space Standardization) manuals commonly used by ESA and other European space agencies foresee that $>15000$ requirements for standard missions (e.g. Euclid) have to be met, mostly on quality standards. This is an expensive and time consuming procedure that greatly reduces the probability of failures, but does not eliminate them. This is demonstrated by the problems that several ESA missions have had to deal with, the most recent ones being on Euclid [^24] https://www.space.com/euclid-guidance-light-sensor-setbacks-commissioning-dark-universe https://europeanspaceflight.com/esa-euclid-telescope-has-an-ice-problem/ and BepiColombo [^25] https://spacenews.com/esa-delays-bepicolombo-orbital-insertion-because-of-thruster-problem/.
@@ 3.1 Metrics for science mission productivity and impact
The high cost of standard space missions should be balanced by a high scientific return. It is not easy to provide an unbiased and quantitative measure of the scientific return of a mission. A standard metric for scientific productivity is the number of refereed papers produced based on a given mission, while a standard metric for evaluating the impact is the number of citations these papers have received. These metrics have the advantage that they are relatively easy to obtain but have important limitations; see Appendix A for a description on data collection and a discussion of their value. Figure 3 shows the productivity and impact of several recent space missions as a function of their mass and cost. The productivity is parameterized as the number of papers per year since operation (YSO) per M\$ of total cost (bus, payload, launch), the impact is parameterized as the Research Impact quotient (Riq) divided by the square root of the total cost in M\$, see Appendix A for definitions.
@@
Of course, simple metrics cannot tell the full story. Large, ambitious missions are more likely to produce transformational results than small, limited missions. Therefore, the real problem is whether space science can take advantage of the benefits of the new space economy also for medium/large missions in addition to nano/small missions. In particular, can ambitious science missions exploit the three pillars of new space?
@@ 4 Toward a new generation of ambitious and affordable space science missions
The development of several cubesat scientific projects in the past few years is a first attempt to take advantage of some of the benefits of the new space economy. The cubesat concept was proposed in 1999 by R. J. Twiggs at Stanford University. The original idea was to provide an educational framework for developing a satellite mission. This implies a short development time, consistent with a master’s degree or PhD courses, and a low cost, affordable by a University. The main limitation at that time was launch cost per kg, which drove the requirement of having spacecraft of just a few kg. After more than 20 years of development, cubesats have moved out of the purely educational sphere to be used for both commercial, scientific and even defence applications. The total cubesat mission cost is in many cases small in absolute value, and thus affordable to Universities and research institutes teams, without the need of large investment by traditional space agencies. As an example, the total cost of the HERMES Pathfinder project, including flight hardware and software development, launch, operation, and science is $\sim$12MEUR, approximately one fourth paid from a European Commission Horizon 2020 grant, and the rest paid from an Italian Ministry of University and Research grant, and from the Agenzia Spaziale Italiana (ASI) funds.
@@
However, while having high impact/dollar (Section 3), scientific cubesats are still rather expensive in terms of cost/kg (section 2.). In particular, the cost of cubesat missions with ambitious requirements like HERMES Pathfinder. The high cost/kg of HERMES Pathfinder has three main contributing factors:
@@
    ambitious scientific and system requirements: temporal resolution of 300-400ns, 7 times better than that of Fermi/GBM; extremely wide energy band from a few keV to a few MeV; sensitivity similar to that of Fermi/GBM; S/C 3 axis stabilized, ADCS system from an higher category (6U); UHF/VHF, S-band and Iridium communications; $>$two year lifetime;
@@
Conversely, Camelot pathfinder missions (GRBAlpha, VZLUSAT2, GRBBeta see[19, 20]), have been developed following an iterative approach, delivering at each stage a minimum viable product, but reaching orbit in a very short time and very low price, still producing relevant scientific results [21, 22].
So far, the developers of scientific cubesat projects have not been able to take advantage of two key elements of cost reduction in the commercial space economy: scale production and vertical integration. The cost of COTS components procured on the market, although much smaller than space-qualified components, remains relatively high. As launch costs continue to decline (Section 2.2) the limitations of mass and volume will be surmountable. The cubesat approach could then be applied to larger spacecraft. We are already witnessing this trend: the number of small satellites predicted to launch in the next decade is declining as typical satellites get heavier [^27] https://spacenews.com/heavier-smallsats-weigh-down-market-forecasts/. The first possible candidates for scientific applications are the smallsats from 50kg to 300kg, that enter the range of the NASA Explorer missions. So far, the realization of the Explorer mission S/C has been contracted to traditional space companies [^28] Ball Aerospace and technology for IXPE, SPHEREx; Nortrop Grumman or Orbital ATK for NuStar, TESS, ICON, COSI; Lockheed Martin for MUSE, respectively. Today, new space companies are exploiting scale production (multi-purpose buses) and vertical integration to offer satellite buses in the 100-1000kg range at low prices. Examples are K2 Space [^29] https://www.k2space.com a company with the goal of providing large satellite buses (P/L up to 1 ton) at low prices ($<$\$15M) and short procurement time (months instead of years or even decades); and APEX303030 https://www.apexspace.com/, with the aim of providing buses to host payloads in the 100-150 kg range at prices of 4-8M \$ and with a procurement time of half a year. Exploiting such standardized buses, without major customization, can make scientific missions cheaper and much more frequent.
Space-based observatories and exploration probes that are developed on timescales of years instead of decades can respond in a much more timely way to emerging scientific problems. More frequent space missions can revolutionize the way space science is done because they can trigger a virtuous loop between opening up new fields of investigation and finding timely responses to the new questions.
@@
(1) a wider funding base. In the old space economy nearly the totality of the basic science missions were paid by governments, and in particular by space agencies, NASA and ESA first among them. And space agencies have their settled ways of working on science missions (see Table 5), which will take time to change, although elements of change have been seen both in the past (the faster, better, cheaper mantra of then NASA Administrator Dan Goldin during the 1990s [21], and in the present (as in the already mentioned case of the re-think of MSR and VIPER).
In contrast,the new space economy is mainly driven by private funding. Private investors usually look for commercial returns on short time scales, while the return from basic science is definitely on a long time scale, if any. So a sudden shift from public to private funding seems unlikely. What can be done by scientists is to work toward a diversification of funding from other entities such as:
@@
    public-private partnership (PPP), with private companies providing advanced technological capabilities as services. PPP was the basis for innovation in space infrastructure, starting with the NASA Commercial Orbital Transportation Services (COTS) program in the 2000s [6, 23]. Other examples of past very fruitful PPPs are PNT and Earth Observation where public funding started or helped develop the systems, and private funding helped develop the market [6];
@@
In the first three cases, technological development starts at TRL 5 or 6. The development of new technologies (TRL 0-5) would remain funded through government funds and, at least partly, from intergovernmental organizations, although philanthropy may contribute. This is of course a key issue because transformational results are often obtained using novel technologies which had to be invented on purpose.
@@
    make the requirements smart. If we want to develop in a systematic way basic science missions in the new space context, we must understand that our scientific requirements must not be carved in stone. Where you end is not where you start. Constraints and requirements can become obsolete, surpassed, and therefore no longer relevant. Sometimes the lesson that ’we did it this way and it worked, so we should do the same again’ is not the right lesson for new circumstances. While choosing ambitious science objectives we must at the same time push for the good, the solid, the robust rather than the best imaginable. Cost should be a major factor in choosing our approach. Once we are able to reach the good on a short timescale, we can improve it at the next step. Of course, descoping is a concept well known to the scientific community and has been applied to many missions in recent decades. Here, though, it is not simply a matter of recognizing at some point that a particular science or system requirement is not achievable at the expected cost or time and eliminating or simplifying it. Rather, it is about planning from the beginning an iterative process to achieve the main science goal, during which the science requirements are continually challenged and questioned, with rapid iteration at each step of the development process. This is quite a radical change with respect to the way a science mission is developed today at NASA and ESA, and to the ways most scientists working on space missions are used to. However, if missions become much more frequent and with much shorter development times, requirements compromises make a lot of sense to speed up the entire process; we can more than recover the science loss in the compromise in a new mission.
@@
    Most academic institutions must cope with rules for the recruitment of personnel and for procurement that are much more cumbersome and stringent than for private companies. Therefore, acceleration of scientific project development times probably requires the setting up of spin-offs or public-private organizations.
The above points imply a major paradigm shift. This shift is unlikely to be made quickly and comprehensively. It is probably more efficient to identify ’case studies’ first. The large programs in Moon exploration of both Western nations and China and its allies, with the Artemis and ILRS programs, offer a path forward. Scientific projects associated with these lunar programs could be placed as a first-case test bed for the changes proposed here. Major resources have been brought into play for lunar programs by both governments and private investors in the hope of creating a truly new lunar space economy. Scientists can ’sell’ the knowledge needed to harvest resources on the Moon, leveraging the ability of the same knowledge to have a high scientific value. In other cases the instruments used for lunar prospecting can be re-tuned to respond to other ambitious scientific questions. Definitely, a lot of good science can be done on or from the Moon [24], and it would be truly bizarre if this science did not benefit from the emerging lunar space economy.

### v2

@@ 2.3 Costs: satellite production
The second major cost of a space mission is the cost related to the satellite procurement, integration, and test. Table 4 reports the cost of several scientific satellites, including payloads, bridging the typical cost of large flagship missions and small explorer missions (1.4-0.42M\$/kg). For comparison, we also report the costs of standard cubesats using COTS components, advanced cubesat, cubesats for commercial constellations, and the SpaceX Starlink satellites.
@@
Satellite Multiplicity Mass kg Cost M\$ Cost M\$/kg
@@
The cost/kg of standard cubesats is 5-20 times lower than the cost/kg of Explorer missions. Of course, standard cubesat are very simple and limited. A better comparison can be made with advanced cubesats, designed specifically for a scientific mission. One good example is HERMES Pathfinder (HERMES-PF), a European constellation of 6 3U cubesat equipped with miniaturized but advanced X-ray and gamma-ray detectors to observe cosmic explosions such as Gamma Ray Bursts and the electromagnetic counterparts of gravitational wave events [19, 20, 21]. The net cost of each HERMES-PF 3U CubeSat is $\sim$\$860.000, including payload (P/L). The spacecraft (S/C) uses selected COTS components, procured through contracts with companies including non-recurring engineering phases to tailor the hardware to the specific application. The P/L and the S/C do not use radhard Electrical, Electronic and Electro-mechanical (EEE) components. HERMES-PF S/C cost/kg is about 145.000 \$/kg (the cost/kg of the P/L is similar), or 2-7 times higher than the standard 3U cubesat but only one-third of that of Explorer missions. The cost of cubesats and mini/midi-sats for commercial constellations is similar or lower to that of standard cubesats, at a level of 20.000-30.000\$/kg  [^23] https://sustainability.e-shape.eu/economics-behind-satellite-megaconstellations/https://www.bloomberg.com/news/features/2017-06-29/the-tiny-satellites-ushering-in-the-new-space-revolution
The cost of SpaceX Starlink satellites (bus and P/L included) is lower than that of cubesats by more than another order of magnitude, reaching an incredibly low value of about 1000\$/kg [^24] https://spacenews.com/starlink-soars-spacexs-satellite-internet-surprises-analysts-with-6-6-billion-revenue-projection/. Interestingly, this cost/kg is about constant for all Starlink satellites despite a weight difference of a factor of 6. This is 20-30 times lower than the cost of a standard cubesat and that of advanced cubesat and mini-sat for Earth observation, more than 100 times lower than the cost of advanced cubesats for science applications, and more than 400 times lower than the cost of Explorer missions.
@@ 3 Public Funding Pros and Cons for Large Missions
So far science missions in space have been paid almost entirely by public funding (governments), and thus they belong to the ’old space’ economy. Table 5 summarizes the pros and cons of this situation.
@@
Figure 2 shows the evolution of NASA and ESA budgets since 2000. NASA’s budget for science, including solar system exploration, Mars exploration, astrophysics, and Earth observation, varied between 25% and 35% of the total between 2000 and 2025. The ESA budget for science decreased from about 15% of the total in 2006-2008 to 8.5% in 2020-2025. This does not include Earth observation, which includes in the ESA system both science and application (the Copernicus program). Figure 2 also shows NASA and ESA science budgets in absolute values, corrected for inflation to 2023. Note the peaks in the NASA budget around 2003-2006 and 2018-2022. The second is mainly due to spending on JWST. The first is due to spending on several smaller missions, including Messenger, MRO, New Horizon, Dawn, Stereo and Fermi, witnessing a period of great diversification in NASA science missions. ESA science budget remained remarkably constant around a value of 600-700 M\$(2023) per year for about 25 years (the 2003 budget was increased by about 100MEUR to take stock of the grounding of Ariane 5, which postponed the launch of Rosetta). ESA budget between 1986, following the agreement of the Horizon 2000 plan, and 2013 is presented by [22]. The budget peaked in 1996 at about M\$(2023), when a negative gradient started because at the 1995 ESA Council Ministerial meeting a flat cash was imposed. The Horizon 2000 plan established Europe as a major player in space science and drove technological innovation in Europe. It was focused on four ”cornerstone” missions, Soho/Cluster, XMM-Newton, Herschel and Rosetta, and several smaller missions including Planck, Mars-Express, Huygens, Integral, ISO, Hipparcos [23]. ESA budget for Earth observation increased consistently from 2008 to 2025 from 500 M\$(2023) to 2700 M\$ (2023), due to the investment of the European Commission in the Copernicus program.
@@
    Use of space qualified parts. As an example, the RAD750 computer that operates in many NASA missions including the Perseverance Mars rover costs several hundred thousands dollars, and is based on 1990 technology with a PowerPC 130-200MHz processor, $\sim$260-400 DMIPS [^25] RAD750® radiation-hardened PowerPC microprocessor BAE Systems https://www.baesystems.com/en-media/uploadFile/20210404045936/1434555668211.pdf. As a comparison, the computer used by the Ingenuity helicopter on board Perseverance is based on the Qualcomm Flight board, which uses a smartphone computer chip, the Qualcomm Snapdragon 801 processor with a 2.5GHz clock, some 20 times more powerful and 20 times cheaper than the RAD750262626https://www.qualcomm.com/news/onq/2021/03/journey-mars-how-our-collaboration-jet-propulsion-laboratory-fostered-innovation. The Snapdragon 801 worked flawlessly for the about five years of cruise and operations of Ingenuity.
@@
    Failure is not an Option leads to large numbers of quality control procedures to reduce the risk of failure. As an example, the ECSS (European Cooperation for Space Standardization) manuals commonly used by ESA and other European space agencies foresee that >15000>15000> 15000 requirements for standard missions (e.g. Euclid) have to be met, mostly on quality standards. This is an expensive and time consuming procedure that greatly reduces the probability of failures, but does not eliminate them. This is demonstrated by the problems that several ESA missions have had to deal with, the most recent ones being on Euclid [^27] https://www.space.com/euclid-guidance-light-sensor-setbacks-commissioning-dark-universe https://europeanspaceflight.com/esa-euclid-telescope-has-an-ice-problem/ and BepiColombo [^28] https://spacenews.com/esa-delays-bepicolombo-orbital-insertion-because-of-thruster-problem/.
@@ 3.1 Metrics for science mission productivity and impact
The high cost of standard space missions should be balanced by a high scientific return. It is not easy to provide an unbiased and quantitative measure of the scientific return of a mission. A standard metric for scientific productivity is the number of refereed papers produced based on a given mission, while a standard metric for evaluating the impact is the number of citations these papers have received. These metrics have the advantage that they are relatively easy to obtain but have important limitations; see Appendix A for a description on data collection and a discussion of their value. Figure 3 shows the productivity and impact of several recent space missions as a function of their mass and cost. Productivity is parameterized as the number of articles per year since operation (YSO) per M\$ of total cost (bus, payload, launch), the impact is parameterized as the Research Impact quotient (Riq) divided by the square root of the total cost in M\$, see Appendix A for definitions.
@@
Of course, simple metrics cannot tell the full story. Large, ambitious missions are more likely to produce transformational results than small, limited missions. Therefore, the real problem is whether space science can take advantage of the benefits of the new space economy also for medium/large missions in addition to nano/micro/small missions. In particular, can ambitious science missions exploit the three pillars of new space?
@@ 4 Toward a new generation of ambitious and affordable space science missions
The development of several cubesat scientific projects in the past few years is a first attempt to take advantage of some of the benefits of the new space economy. The cubesat concept was proposed in 1999 by R. J. Twiggs at Stanford University. The original idea was to provide an educational framework for developing a satellite mission. This implies a short development time, consistent with a master’s or Ph.D. degree courses, and a low cost, affordable by a university. The main limitation at that time was the launch cost per kg, which drove the requirement of having spacecraft of just a few kilograms. After more than 20 years of development, cubesats have moved out of the purely educational sphere to be used for both commercial, scientific, and even defense applications. The total cubesat mission cost is in many cases small in absolute value, and thus affordable to Universities and research institutes teams, without the need of large investment by traditional space agencies. As an example, the total cost of the HERMES Pathfinder project, including flight hardware and software development, launch, operation, and science is $\sim$12MEUR, approximately a fourth paid from a European Commission Horizon 2020 grant, and the rest paid from an Italian Ministry of University and Research grant, and from the Agenzia Spaziale Italiana (ASI) funds.
@@
However, while having high impact/dollar (Section 3), scientific cubesats are still rather expensive in terms of cost/kg (section 2). In particular, the cost of cubesat missions with ambitious requirements like HERMES Pathfinder. The high cost/kg of HERMES Pathfinder has three main contributing factors:
@@
    ambitious scientific and system requirements: temporal resolution of 300-400ns, 5-7 times better than that of Fermi/GBM; extremely wide energy band from a few keV to a few MeV; sensitivity similar to that of Fermi/GBM; S/C 3 axis stabilized, ADCS system from an higher category (6U); UHF/VHF, S-band and Iridium communications; $>$two year lifetime;
@@
In contrast, Camelot pathfinder missions (GRBAlpha, VZLUSAT2, GRBBeta see[26, 27]), have been developed following an iterative approach, delivering at each stage a minimum viable product, but reaching orbit in a very short time and very low price, still producing relevant scientific results [28, 29].
So far, the developers of scientific cubesat projects have not been able to take advantage of two key elements of cost reduction in the commercial space economy: scale production and vertical integration. The cost of COTS components procured on the market, although much smaller than space-qualified components, remains relatively high. As launch costs continue to decline (Section 2.2) the limitations of mass and volume will be surmountable. The cubesat approach could then be applied to larger spacecraft. We are already witnessing this trend: the number of small satellites predicted to launch in the next decade is declining as typical satellites get heavier [^30] https://spacenews.com/heavier-smallsats-weigh-down-market-forecasts/. The first possible candidates for scientific applications are the smallsats from 50kg to 300kg, that enter the range of the NASA Explorer missions. So far, the realization of the Explorer mission S/C has been contracted to traditional space companies [^31] Ball Aerospace and technology for IXPE, SPHEREx; Nortrop Grumman or Orbital ATK for NuStar, TESS, ICON, COSI; Lockheed Martin for MUSE, respectively. Today, new space companies are exploiting scale production (multi-purpose buses) and vertical integration to offer satellite buses in the 100-1000kg range at low prices. Examples are K2 Space [^32] https://www.k2space.com a company with the goal of providing large satellite buses (P/L up to 1 ton) at low prices ($<$\$15M) and short procurement time (months instead of years or even decades); and APEX333333 https://www.apexspace.com/, with the aim of providing buses to host payloads in the 100-150 kg range at prices of 4-8M \$ and with a procurement time of half a year. Exploiting such standardized buses, without major customization, can make scientific missions cheaper and much more frequent.
Space-based observatories and exploration probes that are developed on timescales of years instead of decades can respond much more quickly to emerging scientific problems. More frequent space missions can revolutionize the way space science is done because they can trigger a virtuous loop between opening up new fields of investigation and finding timely responses to the new questions.
@@
(1) a wider funding base. In the old space economy, nearly all basic science missions were paid by governments, and in particular by space agencies, NASA and ESA first among them. And space agencies have their settled ways of working on science missions (see Table 5), which will take time to change, although elements of change have been seen both in the past (the faster, better, cheaper mantra of then NASA Administrator Dan Goldin during the 1990s [21], and in the present (as in the already mentioned case of the rethink of MSR and VIPER).
In contrast,the new space economy is mainly driven by private funding. Private investors usually look for commercial returns on short time scales, while the return from basic science is definitely on a long time scale, if any. So, a sudden shift from public to private funding seems unlikely. What scientists can do is to work toward diversifying funding from other entities such as:
@@
    public-private partnership (PPP), with private companies providing advanced technological capabilities as services. PPP was the basis for innovation in space infrastructure, starting with the NASA Commercial Orbital Transportation Services (COTS) program in the 2000s [7, 30]. Other examples of very fruitful PPPs in the past are PNT and Earth Observation, where public funding started or helped develop the systems, and private funding helped develop the market [7].
@@
In the first three cases, technological development starts with TRL 5 or 6. The development of new technologies (TRL 0-5) would remain funded through government funds and, at least partially, from intergovernmental organizations, although philanthropy may contribute. This is of course a key issue because transformational results are often obtained using novel technologies which had to be invented on purpose.
@@
    make the requirements smart. If we want to develop in a systematic way basic science missions in the new space context, we must understand that our scientific requirements must not be carved in stone. Where you end is not where you start. Constraints and requirements can become obsolete, surpassed, and therefore no longer relevant. Sometimes the lesson that ’we did it this way and it worked, so we should do the same again’ is not the right lesson for new circumstances. While choosing ambitious science objectives, we must at the same time push for the good, the solid, the robust rather than the best imaginable. Cost should be a major factor in choosing our approach. Once we are able to achieve the good in a short timescale, we can improve it in the next step. Of course, descoping is a concept well known to the scientific community and has been applied to many missions in recent decades. Here, though, it is not simply a matter of recognizing at some point that a particular science or system requirement is not achievable at the expected cost or time and eliminating or simplifying it. Rather, it is about planning from the beginning an iterative process to achieve the main science goal, during which the science requirements are continually challenged and questioned, with rapid iteration at each step of the development process. This is quite a radical change with respect to the way a science mission is developed today at NASA and ESA, and to the ways most scientists working on space missions are used to. However, if missions become much more frequent and with much shorter development times, requirements compromises make a lot of sense to speed up the entire process; we can more than recover the science loss in the compromise in a new mission.
@@
    Many academic institutions are not designed to be able to collaborate efficiently with the commercial world. There is a lack of adequate information on business logic in many cases, and a lack of a true risk-oriented culture. We often fall into two extremes: most of the times risk is only a virtual concept that remains on paper; some time however we tend to put all our eggs in one basket. In any case risk and failure are not usually considered positive attributes, and we rarely follow the approach of learning from mistakes. Furthermore, most academic institutions must deal with rules for the recruitment of personnel and for procurement that are much more cumbersome and stringent than for private companies. Therefore, acceleration of scientific project development times probably requires the setting up of spin-offs or public-private organizations.
The above points imply a major paradigm shift. This shift is unlikely to be made quickly and comprehensively. It is probably more efficient to identify case studies first. The large programs in Moon exploration of both Western nations and China and its allies, with the Artemis and ILRS programs, offer a path forward. Scientific projects associated with these lunar programs could be placed as a first-case test bed for the changes proposed here. Major resources have been brought into play for lunar programs by both governments and private investors in the hope of creating a truly new lunar space economy. Indeed, according to a study by the Center for Strategic and International Studies, the size of the lunar economy during the next decade will be of the order of 8-10B\$/yr [^35] https://csis-website-prod.s3.amazonaws.com/s3fs-public/2024-10/241021_Swope_Swimmimg_Upstream_0.pdf.
Scientists can ’sell’ the knowledge needed to harvest resources on the Moon, leveraging the ability of the same knowledge to have a high scientific value. In other cases the instruments used for lunar prospecting can be re-tuned to respond to other ambitious scientific questions. Definitely, a lot of good science can be done on or from the Moon [31, 32], and it would be truly bizarre if this science did not benefit from the emerging lunar space economy. Lunar science and lunar economy can also be good examples of full public-private partnership [33].